### v1

# On Lagrangian formulations for (ir)reducible mixed-antisymmetric higher integer spin fields in Minkowski spaces

Abstract
We extend the results of Lagrangian formulations study to construct gauge-invariant Lagrangians for (ir)reducible integer higher-spin massless and massive representations of the Poincare group with a Young tableau $Y[\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}]$ in $d$-dimensional flat space-time (as the probable candidates to describe the Dark Matter problem beyond the SM). These particles are described within a metric-like formulation by tensor fields with 3 groups of antisymmetric Lorentz indices $\Phi_{\mu^{1}[{\hat{s}_{1}}],\mu^{2}[{\hat{s}_{2}}],\mu^{3}[{\hat{s}_{3}}]}$ on a basis of the BRST method with complete, $Q$, and incomplete, $Q_{c}$, BRST operators. We found unconstrained (with $Q$) and constrained (with $Q_{c}$ and off-shell BRST invariant holonomic constraints) gauge Lagrangian formulations with different configuration spaces and reducibility stages. The deformation procedure to construct interacting gauge model with mixed-antisymmetric fields is proposed.
gauge invariance, BRST approach, cubic vertex, higher spin fields
pacs:
11.30.-j; 11.30.Cp; 11.10.Ef; 11.10.Kk; 11.15. q
1 Introduction
Problems of higher-spin (HS) field theory have been the subject of significant discussions among theoretician community. Interest in HS field theory (for review, see e.g. [1], [2], [3], [4]) is mainly inspired by hope to develop the new Lagrangian models in classical field theory and may be formulate on this ground the new tools for the unification of the fundamental interactions. HS field theory is closely related to (super)string theory [5], [6] which operates with an infinite tower of bosonic and fermionic HS fields in $10$-dimensional space-time with compactified extra dimensions thus generating theirs masses within dimensional reduction mechanism. Because of interacting (super)strings seems to be only unique way to consistently describe evolution of early Universe and involve the interaction of particles with different values of spin, starting from known particles with lower spins (Higgs particle for $s=0$; leptons and quarks for $s=\frac{1}{2}$; photon $W^{\pm},Z$ bosons for $\lambda=1$ and yet undiscovered graviton for $\lambda=2$) the unknown particles both with HS $s>2$ and with generalized spin $s_{(k)}=(s_{1},s_{2},...,s_{k})$ are also admissible. From this perspective such particles may be considered as alternative candidate to describe Dark Matter, see for a review[7], [8] (always surrounding electromagnetically observable particles) beyond the extensions of the Standard Model involving sterile neutrino [9] and dark massive photon [10]. Whereas the problem of LF construction for HS fields in constant curvature spaces have been considered in different approaches (see for review, [11], [12]) both in metric and frame-like formalism for totally, e.g. in [13], [14], [15], [16], [17], [18] and mixed-symmetric irreducible (spin)-tensors fields on $\mathbb{R}^{1,d-1}$ (or AdS space), e.g. in [19], [20], [21], [22], [23], [24], [25], [26], [27] (for $d=4$ as well as for irreducible representations (irreps) with continuous spin, see [28], [29], [30], [31]), the consistent interacting LF with HS fields is still far from its complete realization, in spite for the results (see, e.g., the papers and references therein) for cubic [33], [32], [34], [35], [36], [37], [38] and quartic [39], [40] vertices.
The another way to describe Poincare group ISO(1,d1)ISO(1,d-1) irreps, for $d>4$, is to consider totally and mixed-antisymmetric (spin)-tensors fields, for which the LF for the former have been found in [41], [42], [43] and for massless HS fields with antisymmetric spin ${s}{[2]}=(\hat{s}_{1},\hat{s}_{2})$ in [44], [45], [46], [47] and in BRST (Becchi-Rouet-Stora-Tyutin) approaches with complete and incomplete BRST operators for massive and massless cases in [48].
[…]
$Q_{c}=C^{A}o_{A}+\frac{1}{2}{C}^{A}{C}^{B}f^{D}_{BA}{\mathcal{P}}_{D}(-1)^{\epsilon({p}_{D})+\epsilon({o}_{A})},\ \ Q=C^{I}o_{I}+\frac{1}{2}{C}^{I}{C}^{J}f^{K}_{JI}{\mathcal{P}}_{K}(-1)^{\epsilon({o}_{K})+\epsilon({o}_{I})}$ (1)
with structural constants $f^{K}_{JI}(f^{D}_{BA})$ from involution relations: $o^{I},o_{J}\}=f^{K}_{IJ}o_{K}$, with operators of ghost coordinates $C^{I}$, momenta ${\mathcal{P}}_{J}$ of opposite to $o_{I}$ Grassmann parities and with $C^{I},{\mathcal{P}}_{J}\}=\delta^{I}_{J}$.
The first purpose of the paper is to construct general gauge-invariant Lagrangians for both massless and massive mixed-symmetry tensor fields of rank $s_{1}+s_{2}+s_{3}$, with any integer numbers $s_{1}\geq s_{2}\geq s_{3}\geq 1$ for $s_{1}<[d/2]$ in $\mathbb{R}^{1.d-1}$, as elements of Poincare-group ISO(1,d1)ISO(1,d-1) irreps with a Young tableau having $3$ columns. Another aim is to formulate the deformation procedure to construct interacting gauge model with mixed-antisymmetric fields.
In the Section 2 we firstly construct LF with complete BRST operator for HS field. A LF with incomplete BRST operators and appropriate set of holonomic constraints is derived in Section 3. Section 4 contains the concept of deformation procedure for interacting LF of $p$ samples of irreducible HS fields subject to $Y[\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}]$.
We use the definitions and notations from [36], [37], [48] for a metric tensor $\eta_{\mu\nu}$ and $(\epsilon,gh_{H})(F)$, [,}[\ ,\ \}, $[x]$, s[3]{s[3]}, $({s}{[3]})_{p}$, $A^{[i}B^{j]k}$ for the values of Grassmann parity and ghost number of a quantity $F$, supercommutator, the integer part of real $x$, for the antisymmetric spin with components $(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3})$, for the $p$-multiple $({s}^{1}_{[3]},{s}^{2}_{[3]},...,{s}^{p}_{[3]})$ and for antisymmetrization $A^{[i}B^{j]k}=A^{i}B^{jk}-A^{j}B^{ik}$.
2 Lagrangian formulation with complete BRST operator \\
Here, we shortly recall [48] that a massless (massive) integer spin irreducible representation of Poincare group in a Minkowski space $\mathbb{R}^{1,d-1}$ which is described by a real-valued tensor field $\Phi_{\mu^{1}[{\hat{s}_{1}}],\mu^{2}[{\hat{s}_{2}}],\mu^{3}[{\hat{s}_{3}}]}\hskip-1.99997pt$ $\equiv$ $\hskip-1.99997pt\Phi_{\mu^{1}_{1}\ldots\mu^{1}_{\hat{s}_{1}},...,\mu^{3}_{1}\ldots\mu^{3}_{\hat{s}_{3}}}$ of rank $\sum_{i}^{3}\hat{s}_{i}$ and generalized spin $\mathbf{s}=(3,3,...,3,2,2,...,2,1$ ,,1),...,1), ($\hat{s}_{1}\geq\hat{s}_{2}\geq s_{3}>0,\hat{s}_{1}\leq[d/2])$ to be corresponding to a Young tableau with $3$ columns of height $\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}$, respectively (with omitting the symbol $\hat{\phantom{s}}"$ under $\hat{s}_{i}$)
[…]
This field is antisymmetric with respect to the permutations of each type of Lorentz indices $\mu^{i}$, and obeys to the wave (Klein-Gordon) (3), divergentless (4), traceless (5) and mixed-antisymmetry equations (6) :
$(\partial^{\mu}\partial_{\mu}(+m^{2}))\Phi_{\mu^{1}{[s_{1}]},\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}=0,$ (3)
$(\partial_{i}\Phi)_{{\mu}^{1}[{s_{1}-\delta_{i1}}],...,{\mu}^{3}[{s_{3}-\delta_{i3}}]}\equiv\partial^{\mu^{i}_{l_{i}}}\Phi_{\mu^{1}{[s_{1}]},\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}=0,\texttt{ for }1\leq l_{i}\leq s_{i},\ i=1,2,3,$ (4)
$(\texttt{Tr}^{ij}\Phi)_{\mu^{1}{[s_{1}]},\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}\equiv\eta^{\mu^{i}_{l_{i}}\mu^{j}_{l_{j}}}\Phi_{\mu^{1}{[s_{1}]},\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}=0,\ \ 1\leq i<j\leq 3,$ (5)
$(Y^{ij}\Phi)_{\mu^{1}[{s_{1}}],...,[\mu^{i}[{s_{i}}],...,\mu^{j}_{l_{j}}],\hat{\mu}^{j}[{s_{j}-1}],...,\mu^{3}[{s_{3}}]}\equiv-(-1)^{l_{j}}\Phi_{[...,[\mu^{i}[{s_{i}}],...,\mu^{j}_{l_{j}}]\underbrace{\mu^{j}_{1}...\mu^{j}_{l_{j}-1}}\mu^{j}_{l_{j}}]...\mu^{j}_{s_{j}},...,\mu^{3}[{s_{3}}]}=0,$ (6)
with use the notations for the operations of divergences $(\partial_{i}\Phi)$, mixed traces $(\texttt{Tr}^{ij}\Phi)$, Young-antisymmetrizations $(Y^{ij}\Phi)$. The bracket below in (6) denote that the indices in it do not include in antisymmetrization.
It is easy to see, the system (3)–(6) is equivalent to set of primary constraints on the string=like vector $|\Phi\rangle\in\mathcal{H}^{f}$ in an auxiliary Fock space $\mathcal{H}^{f}$, generated by $3$ pairs of Grassmann-odd (antisymmetric basis) oscillators $\hat{a}^{i+}_{\mu^{i}}(x)$, $\hat{a}^{j}_{\nu^{j}}(x)$, $\mu^{i},\nu^{j}=0,1...,d-1$:
$\hskip-8.61108pt|\Phi\rangle=\sum_{s_{1}=0}^{[d/2]}\sum_{s_{2}=0}^{s_{1}}\sum_{s_{3}=0}^{s_{2}}\frac{\imath^{\sum_{p=1}^{3}s_{p}}}{{s_{1}!s_{2}!s_{3}!}}\Phi_{\mu^{1}{[s_{1}]},\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}\,\prod_{i=1}^{3}\prod_{l_{i}=1}^{s_{i}}\hat{a}^{+\mu^{i}_{l_{i}}}_{i}|0\rangle,\ \ \{\hat{a}^{i}_{\mu^{i}},\hat{a}_{\nu^{j}}^{j+}\}=-\eta_{\mu^{i}\nu^{i}}\delta^{ij},$ (7)
$\hskip-8.61108pt\bigl{(}{l}_{0},\,{l}^{i},\,l^{ij},\,t^{ij},\,g_{0}^{i}-(s_{i}-\textstyle\frac{d}{2})\bigr{)}|\Phi\rangle=\bigl{(}\partial^{\mu}\partial_{\mu},\,-i\hat{a}^{i}_{\mu}\partial^{\mu},\,\textstyle\frac{1}{2}\hat{a}^{i}_{\mu}\hat{a}^{j\mu},\,\hat{a}^{i+}_{\mu}\hat{a}^{j\mu},\,-\hat{a}^{i+}_{\mu}\hat{a}^{\mu{}i}-\frac{d}{2}\bigr{)}|\Phi\rangle=\vec{0}.$ (8)
The set of operators $\{o_{\alpha}\}$ = $\bigl{\{}{{l}}_{0},{l}^{i},l^{ij},t^{ij}\bigr{\}}$ (for $l_{0}$ corresponding to hamiltonian $H_{0}$ from BFV method [51]), which generate $(1+6)$ even and $3$ odd, ${l}^{i}$, primary constraints (8), and number particles operators, $g_{0}^{i}$ should form closed superalgebra with respect to [,}[\ ,\ \}-multiplication and hermitian conjugation with inner product
$\langle{\Psi}|\Phi\rangle$ $=$ $\int d^{d}x\sum_{s_{1}=0}^{[d/2]}\sum_{s_{2}=0}^{s_{1}}\sum_{s_{3}=0}^{s_{2}}\frac{(-1)^{\sum_{p}s_{p}}}{s_{1}!s_{2}!s_{3}!}\Psi^{\star}_{\mu^{1}[{s_{1}}],\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}(x)\Phi^{\mu^{1}[{s_{1}}],\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}(x).$ (9)
As the result, the set of $\{o_{\alpha}\}$ extended by means of the operators,
$\hskip-8.61108pt\bigl{(}{l}^{i+},\ l^{ij+},\ t^{ij+}\bigr{)}=\bigl{(}-i\hat{a}^{i+}_{\mu}\partial^{\mu},\ \textstyle\frac{1}{2}\hat{a}^{j+}_{\mu}\hat{a}^{i\mu+},\ \hat{a}^{j+}_{\mu}\hat{a}^{i\mu}\bigr{)},\$(10)
forms superalgebra Lie of operators $\{o_{I}\}=$ $\{o_{A};o_{a},o_{a}^{+};g_{0}^{i}\}$, for isometry subsuperalgebra of $\{o_{A}\}=$ $\{{l}_{0},{l}^{i},{l}^{i+}\}$ and subalgebra $\{o_{a},o_{a}^{+};g_{0}^{i}\}$ with operators of second-class (holonomic) constraints $\{o_{a},o_{a}^{+}\}$ for $o_{a}=(l^{ij},t^{ij})$.The multiplication law for $o_{I}$ is given by the compact relations
$o_{I},\ o_{J}]=f^{K}_{IJ}o_{K}:\ [o_{a},\;o_{b}^{+}\}=f^{c}_{ab}o_{c}+\Delta_{ab}(g_{0}^{i}),\ f^{K}_{IJ}=-(-1)^{\varepsilon(o_{I})\varepsilon(o_{J})}f^{K}_{JI},$ (11)
with invertible matrix $\|\Delta_{ab}(g_{0}^{i})\|$ in $\mathcal{H}^{f}$ on the solutions for the system $(o_{a},o_{A})|\Phi\rangle=0$. Explicitly, the structure constants $f^{K}_{IJ}$ determined from the multiplication table 1.
$\downarrow$$\rightarrow$] $t^{\prime}_{12}$ $t^{\prime}_{13}$ $t^{\prime}_{23}$ $t^{+\prime}_{12}$ $t^{+\prime}_{13}$ $t^{+\prime}_{23}$ $l^{+\prime}_{12}$ $l^{+\prime}_{13}$ $l^{+\prime}_{23}$ $g^{\prime}_{01}$ $g^{\prime}_{02}$ $g^{\prime}_{03}$
$t^{\prime}_{12}$ 0 0 $t^{\prime}_{13}$ $g^{\prime}_{01}-g^{\prime}_{02}$ $t^{+\prime}_{23}$ 0 0 0 $l^{+\prime}_{13}$ $t^{\prime}_{12}$ $t^{\prime}_{12}$ 0
$t^{\prime}_{13}$ 0 0 0 $t^{\prime}_{23}$ $g^{\prime}_{01}-g^{\prime}_{03}$ $t^{\prime}_{12}$ 0 0 $l^{+\prime}_{12}$ $t^{\prime}_{13}$ 0 $t^{\prime}_{13}$
$t^{\prime}_{23}$ $t^{\prime}_{13}$ 0 0 0 $t^{+\prime}_{12}$ $g^{\prime}_{02}-g^{\prime}_{03}$ 0 $l^{+\prime}_{12}$ 0 0 $t^{\prime}_{23}$ $t^{\prime}_{23}$
$t^{+\prime}_{12}$ $g^{\prime}_{01}+g^{\prime}_{02}$ $t^{\prime}_{23}$ 0 0 0 $t^{+\prime}_{13}$ 0 $l^{+\prime}_{23}$ 0 $t^{+\prime}_{12}$ $t^{+\prime}_{12}$ 0
$t^{+\prime}_{13}$ $t^{+\prime}_{23}$ $g^{\prime}_{01}+g^{\prime}_{03}$ $t^{+\prime}_{12}$ 0 0 0 $l^{+\prime}_{23}$ 0 0 $t^{+\prime}_{13}$ 0 $t^{+\prime}_{13}$
$t^{+\prime}_{23}$ 0 $t^{\prime}_{12}$ $g^{\prime}_{02}+g^{\prime}_{03}$ $t^{+\prime}_{13}$ 0 0 $l^{+\prime}_{13}$ 0 0 0 $t^{+\prime}_{23}$ $t^{+\prime}_{23}$
$l^{\prime}_{1}$ $l^{\prime}_{2}$ $l^{\prime}_{3}$ 0 0 0 0 $\frac{l^{+\prime}_{2}}{2}$ $\frac{l^{+\prime}_{3}}{2}$ 0 $l^{\prime}_{1}$ 0 0
$l^{\prime}_{2}$ 0 0 $l^{\prime}_{3}$ $l^{\prime}_{1}$ 0 0 $\frac{l^{+\prime}_{1}}{-2}$ 0 0 0 $l^{\prime}_{2}$ 0
$l^{\prime}_{3}$ 0 0 0 0 $l^{\prime}_{1}$ $l^{\prime}_{2}$ 0 $\frac{l^{+\prime}_{1}}{-2}$ $\frac{l^{+\prime}_{2}}{-2}$ 0 0 $l^{\prime}_{3}$
$l^{+\prime}_{1}$ 0 0 0 $l^{+\prime}_{2}$ $l^{+\prime}_{3}$ 0 0 0 0 $l^{+\prime}_{1}$ 0 0
$l^{+\prime}_{2}$ $l^{+\prime}_{1}$ 0 0 0 0 $l^{+\prime}_{3}$ 0 0 0 0 $l^{+\prime}_{2}$ 0
$l^{+\prime}_{3}$ 0 $l^{+\prime}_{1}$ $l^{+\prime}_{2}$ 0 0 0 0 0 0 0 0 $l^{+\prime}_{3}$
$l^{\prime}_{12}$ 0 $l^{\prime}_{23}$ $l^{\prime}_{13}$ 0 0 0 $\frac{g^{\prime}_{01}+g^{\prime}_{02}}{-4}$ $\frac{t^{+\prime}_{23}}{4}$ $\frac{t^{+\prime}_{13}}{-4}$ $l^{\prime}_{12}$ $l^{\prime}_{12}$ 0
$l^{\prime}_{13}$ $l^{\prime}_{23}$ 0 0 0 0 $l^{\prime}_{12}$ $\frac{t^{\prime}_{23}}{4}$ $\frac{g^{\prime}_{01}+g^{\prime}_{03}}{-4}$ $\frac{t^{+\prime}_{12}}{4}$ $l^{\prime}_{13}$ 0 $l^{\prime}_{13}$
$l^{\prime}_{23}$ 0 0 0 $l^{\prime}_{13}$ $l^{\prime}_{12}$ 0 $\frac{t^{\prime}_{13}}{4}$ $\frac{t^{\prime}_{12}}{4}$ $\frac{g^{\prime}_{02}+g^{\prime}_{03}}{-4}$ 0 $l^{\prime}_{23}$ $l^{\prime}_{23}$
$l^{+\prime}_{12}$ 0 0 0 0 $l^{+\prime}_{23}$ $l^{+\prime}_{13}$ 0 0 0 $l^{+\prime}_{12}$ $l^{+\prime}_{12}$ 0
$l^{+\prime}_{13}$ 0 0 $l^{+\prime}_{12}$ $l^{+\prime}_{23}$ 0 0 0 0 0 $l^{+\prime}_{13}$ 0 $l^{+\prime}_{13}$
$l^{+\prime}_{23}$ $l^{+\prime}_{13}$ $l^{+\prime}_{12}$ 0 0 0 0 0 0 0 0 $l^{+\prime}_{23}$ $l^{+\prime}_{23}$
$g^{\prime}_{01}$ $t^{\prime}_{12}$ $t^{\prime}_{13}$ 0 $t^{+\prime}_{12}$ $t^{+\prime}_{13}$ 0 $l^{+\prime}_{12}$ $l^{+\prime}_{13}$ 0 0 0 0
$g^{\prime}_{02}$ $t^{\prime}_{12}$ 0 $t^{\prime}_{23}$ $t^{+\prime}_{12}$ 0 $t^{+\prime}_{23}$ $l^{+\prime}_{12}$ 0 $l^{+\prime}_{23}$ 0 0 0
$g^{\prime}_{03}$ 0 $t^{\prime}_{13}$ $t^{\prime}_{23}$ 0 $t^{+\prime}_{13}$ $t^{+\prime}_{23}$ 0 $l^{+\prime}_{13}$ $l^{+\prime}_{23}$ 0 0 0
Table 1: HS symmetry superalgebra $\mathcal{A}(Y[3],\mathbb{R}^{1,d-1})$.
Note the indices $i,j,k$ are raising and lowering by Euclidian metric tensors $\delta^{ij}$, $\delta_{ij}$, $\delta^{i}_{j}$
The algebra of the operators $o_{I}$ is called the integer HS symmetry superalgebra in Minkowski space with a Young tableau having $3$ columns and denoted as $\mathcal{A}(Y[3],\mathbb{R}^{1,d-1})$.
The subalgebra with holonomic constraints $(o_{a},o_{a}^{+})$ is isomorphic to algebra $su(4)$. Its presence means for the construction of LF with complete BRST operator a necessity of conversion into a system of deformed first-class operator constraints $(o_{a},o_{a}^{+})\to(O_{a},O_{a}^{+})$. Within additive conversion receipt, see for instance, [26], [42], [59], which implies the enlarging of $o_{I}$ to $O_{I}=o_{I}+o^{\prime}_{I}$, where additional parts $o^{\prime}_{I}$ are given on a new Fock space $\mathcal{H}^{\prime}$ being independent on $\mathcal{H}^{f}$ in a sense that $\mathcal{H}^{\prime}\bigcap\mathcal{H}^{f}=\emptyset$. The operators $O_{I},o^{\prime}_{I}$ satisfy the same algebra as one (11) for $o_{I}$. From Verma module construction for $su(4)$ with basis $\{|\vec{N}\rangle_{V}\}$
$|\vec{N}\rangle_{V}\ \stackrel{{\scriptstyle def}}{{=}}\ |\vec{n}_{ij},\vec{p}_{rs}\rangle_{V}\ =\ \prod_{i=1}^{2}\Bigr{[}\prod_{j=i+1}^{3}\bigl{(}l^{\prime+}_{ij}\bigr{)}{}^{n_{ij}}\Bigl{]}\prod_{r=1}^{2}\Bigr{[}\prod_{s=r+1}^{3}\bigl{(}t^{\prime+}_{rs}\bigr{)}{}^{p_{rs}}\Bigl{]}|0\rangle_{V},\$(12)
$\texttt{for}\ \big{(}l^{\prime}_{ij},t^{\prime}_{rs};g_{0}^{i\prime}\big{)}|0\rangle_{V}=\big{(}\vec{0}_{ij},\vec{0}_{rs};h^{i}\big{)}|0\rangle_{V},\ \ {}_{V}\langle 0|0\rangle_{V}=1$
(with some $\mathbb{R}$-valued numbers $h^{i}$) it follows from the mapping between basis $\{|\vec{N}\rangle_{V}\}$ (12) and the one $\{|\vec{N}\rangle\}\equiv$ $\{\prod_{i=1}^{2}\prod_{j=i+1}^{3}\bigl{(}b^{+}_{ij}\bigr{)}{}^{n_{ij}}\prod_{r=1}^{2}\prod_{s=r+1}^{3}\bigl{(}d^{+}_{rs}\bigr{)}{}^{p_{rs}}|0\rangle\}$ in new Fock space $\mathcal{H}^{\prime}$, the scalar oscillator representation for $o^{\prime}_{I}$
$g^{1\prime}_{0}$ $=$ $b_{12}^{+}b_{12}+b_{13}^{+}b_{13}-d_{12}^{+}d_{12}-d_{13}^{+}d_{13}+h^{1}$
$g^{2\prime}_{0}$ $=$ $b_{12}^{+}b_{12}+b_{23}^{+}b_{23}-d_{23}^{+}d_{23}+d_{12}^{+}d_{12}+h^{2}$ (13)
$g^{3\prime}_{0}$ $=$ $b_{13}^{+}b_{13}+b_{23}^{+}b_{23}+d_{13}^{+}d_{13}+d_{23}^{+}d_{23}+h^{3}\,$,
$\big{(}l^{\prime+}_{ij},\,t^{+\prime}_{12},\,t^{+\prime}_{13},\,t^{+\prime}_{23}\big{)}$ $=$ $\big{(}b_{ij}^{+},\,d_{12}^{+}-b_{23}^{+}b_{13},\,d_{13}^{+}+b_{23}^{+}b_{12},\,d_{23}^{+}-d_{13}^{+}d_{12}-b_{13}^{+}b_{12}\big{)}$ (14)
for the elements $l^{\prime}_{lm}$ for $l<m$
$l^{\prime}_{12}$ $=$ $\bigl{(}-h^{1}-h^{2}-b_{12}^{+}b_{12}-b_{13}^{+}b_{13}-b_{23}^{+}b_{23}+d_{13}^{+}d_{13}\bigr{)}b_{12}-b_{13}b_{12}b^{+}_{13}+b_{13}d^{+}_{23}-b_{23}d^{+}_{13}$ (15)
$l^{\prime}_{13}$ $=$ $\bigl{(}-h^{1}-h^{3}+b_{12}^{+}b_{12}-b_{13}^{+}b_{13}-b_{23}^{+}b_{23}+d_{12}^{+}d_{12}-d_{23}^{+}d_{23}\bigr{)}b_{13}$
$-b_{12}b_{13}+\bigl{(}-h^{2}+h^{3}+d_{23}^{+}d_{23}\bigr{)}b_{23}+b_{23}d^{+}_{12}$
$l^{\prime}_{23}$ $=$ $-\bigl{(}h^{2}+h^{3}+b_{12}^{+}b_{12}+b_{13}^{+}b_{13}+b_{23}^{+}b_{23}+d_{12}^{+}d_{12}+d_{13}^{+}d_{13}\bigr{)}b_{23}+b_{13}d_{13}d^{+}_{23}$
$-\bigl{(}h^{1}-h^{3}-d_{12}^{+}d_{12}-d_{13}^{+}d_{13}-d_{23}^{+}d_{23}\bigr{)}b_{12}d_{13}$
$+\bigl{(}h^{1}-h^{2}-d_{12}^{+}d_{12}-d_{13}^{+}d_{13}+d_{23}^{+}d_{23}\bigr{)}b_{13}d_{12}-\bigl{(}h^{2}-h^{3}-d_{23}^{+}d_{23}\bigr{)}b_{12}d_{12}d_{23}$
and for $t^{\prime}_{lm}$,
$t^{\prime}_{12}$ $=$ $-b_{13}^{+}b_{23}+\bigl{(}h^{1}-h^{2}-d_{13}^{+}d_{13}+d_{23}^{+}d_{23}-d_{12}^{+}d_{12}\bigr{)}d_{12}+d_{23}^{+}d_{13},$ (18)
$t^{\prime}_{13}$ $=$ $b_{12}^{+}b_{23}+\bigl{(}h^{1}-h^{3}-d_{12}^{+}d_{12}-d_{13}^{+}d_{13}-d_{23}^{+}d_{23}\bigr{)}d_{13}+\bigl{(}h^{2}-h^{3}-d_{23}^{+}d_{23}\bigr{)}d_{12}d_{23},$ (19)
$t^{\prime}_{23}$ $=$ $-b_{12}^{+}b_{13}+\bigl{(}h^{2}-h^{3}-d_{23}^{+}d_{23}\bigr{)}d_{23}-d_{12}^{+}d_{13}.$ (20)
The operators $o^{\prime}_{I}$ are hermitian with respect to modified (by operator $K^{\prime}$ intertwining $\{|\vec{N}\rangle_{V}\}$ and $\{|\vec{N}\rangle\}$ (32)) scalar product. Having in mind the rules of the dimensional reduction from massless theory in $\mathbb{R}^{1,d}$ space-time to massive one in $\mathbb{R}^{1,d-1}$ with fermionic oscillators $f_{i}^{+},f_{i}$,
$\partial^{M}=(\partial^{\mu},-\imath m)\,$, $\hat{a}^{M}_{i}=(\hat{a}^{\mu}_{i},f_{i})\,$, $\hat{a}^{M{}+}_{i}=(\hat{a}^{\mu{}+}_{i},f_{i}^{+})\,$, (21)
$M=0,1,\ldots,d\,$, $\mu=0,1,\ldots,d-1\,$, $\eta^{MN}=diag(1,-1,\ldots,-1,-1)\,$, (22)
the converted operators $O_{I}$ now for massive case are derived as well.
2.1 BRST operator and Lagrangian formulation
Complete BRST operator $Q$ is included into nilpotent operator $Q^{\prime}$, (due to number particle operators $G_{0}^{i}$ presence in $O_{I}$) and may be easily found according to prescription (1). $Q^{\prime}$ can be presented via incomplete BRST operator $Q_{c}$ with only differential first-class constraints $\{l_{0},l_{i},l_{i}^{+}\}$ and with complete BRST-extended converted second-class constraints $\{\mathcal{O}_{a}.\mathcal{O}^{+}_{a}\}$ as
$Q^{\prime}$ $=$ $Q_{c}+\sum_{i<j}\Big{(}\eta_{ij}\mathcal{L}{}^{+}_{ij}+\mathcal{L}_{ij}\eta^{+}_{ij}+\mathcal{T}_{ij}\vartheta_{ij}^{+}+\vartheta_{ij}\mathcal{T}^{+}_{ij}\Big{)}+\sum_{i}[\eta^{G}_{i}\sigma^{i}+\imath\mathcal{B}^{i}\mathcal{P}_{i}^{G}],$ (23)
with definite operators $\mathcal{B}^{i}$, generalized spin operator $\sigma^{i}$ and with complete traceless and Young constraints (and theirs hermitian conjugated ones)
$\sigma^{i}\hskip-1.99997pt$ $=$ $\hskip-1.99997ptG_{0i}-q_{i}^{+}p_{i}-q_{i}p_{i}^{+}+\sum_{j\neq i}[\eta_{ij}^{+}\mathcal{P}_{ij}-\eta_{ij}\mathcal{P}^{+}_{ij}]-\sum_{j>i}[\vartheta^{+}_{ij}\lambda_{ij}-\vartheta_{ij}\lambda^{+}_{ik}]+\sum_{j<i}[\vartheta^{+}_{ji}\lambda_{ji}-\vartheta_{ji}\lambda^{+}_{ji}$], (24)
$\mathcal{L}_{ij}\hskip-1.99997pt$ $=$ $\hskip-1.99997pt\widehat{L}^{\prime}_{ij}+\hskip-1.49994pt\sum_{p<j}\hskip-1.99997pt\vartheta_{ip}^{+}\mathcal{P}_{pj}+\hskip-1.99997pt\sum_{j<p\leq 3}\hskip-3.00003pt\vartheta_{jp}^{+}\mathcal{P}_{ip}-\hskip-1.99997pt\sum_{j<p\leq 3}\hskip-3.00003pt\vartheta_{ip}^{+}\mathcal{P}_{jp}+\hskip-1.00006pt\frac{1}{4}\Big{\{}\hskip-1.00006pt\sum_{3\geq p>j}\hskip-3.00003pt\big{[}\eta_{ip}\lambda^{+}_{jp}-\hskip-1.00006pt\eta_{jp}\lambda^{+}_{ip}\big{]}+\sum_{p<j}\hskip-1.99997pt\eta_{pj}\lambda^{+}_{ip}\Big{\}}$, (25)
$\mathcal{T}_{ij}$ $=$ $\widehat{T}^{\prime}_{ij}+\sum_{p>j}\vartheta^{+}_{jp}\lambda_{ip}-\sum_{3\geq p<j}[\vartheta_{ip}\lambda_{pj}-\vartheta_{pj}\lambda_{ip}]+\sum_{3\geq p>j}\eta_{jp}\mathcal{P}_{ip}^{+}-\sum_{i<p<j}\eta_{pj}\mathcal{P}_{ip}^{+}+\sum_{p<i}\eta_{pj}\mathcal{P}_{pi}^{+},$ (27)
$\texttt{for}\ \big{(}\widehat{L}^{\prime}_{ij},\widehat{T}^{\prime}_{ij}\big{)}\ =\ \big{(}L_{ij}+\frac{1}{2}q_{[i}p_{j]},\,{T}_{ij}+(q_{j}p_{i}^{+}+q_{i}^{+}p_{j})\big{)}$
to be incomplete BRST-extended operator constraints $\big{(}\widehat{L}^{\prime}_{ij},\widehat{T}^{\prime}_{ij}\big{)}$, Sum of $Q_{c}$ with rest term without number particles ghost coordinates, $\eta^{G}_{i}$, and momenta, $\mathcal{P}_{i}^{G}$ determines complete BRST operator $Q$ for only converted constraint system with extracted zero-mode ghosts
$Q$ $=$ $Q_{c}+\sum_{i<j\leq 3}\big{(}\mathcal{L}_{ij}\eta^{+}_{ij}+\mathcal{T}_{ij}\vartheta_{ij}^{+}+h.c,\big{)},\ \ (\epsilon,gh_{H})Q=(1,1).$ (28)
(for $gh_{H}(\mathcal{C}^{I})=-gh_{H}(\mathcal{P}_{I})=1$). Here, the Grassman-even ghost coordinates $q_{i}^{+}$, $q_{i}$ momenta $p_{j}$, $p_{j}^{+}$ for Grassman-odd basis elements $l_{i}$, $l_{i}^{+}$ of the superalgebra $\mathcal{A}_{C}(Y[k],\mathbb{R}^{1,d-1})$, respectively, and Grassman-odd ghost coordinates $\eta_{0}$, $\eta_{ij}^{+}$, $\eta_{ij}$, $\vartheta_{rs}^{+}$, $\vartheta_{rs}$, $\eta_{i}^{G}$ and momenta $\mathcal{P}_{0}$, $\mathcal{P}_{ij}$, $\mathcal{P}_{ij}^{+}$, $\lambda_{rs}$, $\lambda_{rs}^{+}$, $P^{G}_{i}$ for Grassman-even basis elements $l_{0}$, $L_{ij}$, $L_{ij}^{+}$, $T_{rs}$, $T_{rs}^{+}$, $G_{0}^{i}$ of $\mathcal{A}_{C}(Y[k],\mathbb{R}^{1,d-1})$, respectively, are introduced subject to non-trivial commutation relations,
$[q_{i}^{+},p_{j}]=[p_{i}^{+},q_{j}]=\delta_{ij},\qquad\{\eta_{0},\mathcal{P}_{0}\}=\imath,\qquad\{\eta^{G}_{i},\mathcal{P}^{G}_{j}\}=\imath\delta_{ij},$ (29)
$\{\eta_{ij}^{+},\mathcal{P}_{i^{\prime}j^{\prime}}\}=\{\eta_{ij},\mathcal{P}_{i^{\prime}j^{\prime}}^{+}\}=\delta_{ij,i^{\prime}j^{\prime}}\{\vartheta_{rs},\lambda_{r^{\prime}s^{\prime}}^{+}\}=\{\vartheta_{rs}^{+},\lambda_{r^{\prime}s^{\prime}}\}=\delta_{rs,r^{\prime}s^{\prime}}.$ (30)
The hermitian conjugation is understood in the sense of the rule
$Q^{\prime+}K\ =\ KQ^{\prime},\ (\mathcal{L}_{ij},\mathcal{T}_{ij})^{+}K\ =\ K(\mathcal{L}^{+}_{ij},\mathcal{T}^{+}_{ij})\texttt{ with }K\ =\ 1\otimes K^{\prime}\otimes 1_{gh}.$ (31)
with non-degenerate Grassmann-even operator $K$ to be determined on the total Hilbert space, $\mathcal{H}_{tot}=$ $\mathcal{H}^{f}\otimes\mathcal{H}^{\prime}$ $\otimes\mathcal{H}_{gh}$. It can be constructed from the whole set of oscillators, whereas $K^{\prime}$ from only $b_{ij}^{(+)}$ and $d_{rs}^{(+)}$:
$K^{\prime}$ $=$ $\hskip-3.00003pt\sum_{n_{ij},p_{rs},n^{\prime}_{ij},n^{\prime}_{rs}=0}^{\infty}\prod_{I,r=1}^{2}\prod_{j\geq i,s\geq r}^{3}\hskip-1.99997pt\frac{{b}_{1j}^{+n_{ij}}}{(n_{ij})!}\frac{{d}_{rs}^{+p_{rs}}}{(p_{rs})!}|0\rangle{}_{V}\langle\vec{n}_{ij},\vec{p}_{rs}|\vec{n}{}^{\prime}_{ij},\vec{p}{}^{\prime}_{rs}\rangle_{V}\langle 0|\hskip-1.99997pt\prod_{I,r=1}^{2}\prod_{j\geq i,s\geq r}^{3}\hskip-3.00003pt\frac{{b}_{1j}^{n^{\prime}_{ij}}}{(n^{\prime}_{ij})!}\frac{{d}_{rs}^{p^{\prime}_{rs}}}{(p^{\prime}_{rs})!},$ (32)
Thus, we have constructed a Hermitian complete BRST operator for the HS symmetry superalgebra $\mathcal{A}_{C}(Y[3],\mathbb{R}^{1,d-1})$ of converted operators $O_{I}$. Including oscillators $f_{i}^{+}$ and the ghost operators, we extend our basic vector $|\Phi\rangle$ (7) given in $\mathcal{H}^{f}$ to
$|\chi\rangle$ $=$ $\sum_{\{n\}_{b}=0}^{\infty}\sum_{\{n\}_{f}=0}^{1}\eta_{0}^{n_{\eta_{0}}}\prod_{l=1}^{3}(\eta_{l}^{G})^{n_{l}}q_{l}^{+n_{q_{l}}}p_{l}^{+n_{p_{l}}}f_{l}^{+n_{f_{l}}}\prod_{1\leq i<j\leq 3}\eta_{ij}^{+n_{\eta_{ij}}}\mathcal{P}_{ij}^{+n_{P_{ij}}}{b}_{1j}^{+n_{ij}}$
[…]
where $\{n\}_{b}=n_{q_{l}},n_{p_{l}},n_{ij},p_{rs}$ and $\{n\}_{f}=n_{\eta_{0}},n_{\eta_{ij}},n_{P_{ij}},n_{\vartheta_{rs}},n_{\lambda_{rs}},n_{f_{l}}$
From the BRST-like complex, with equation determining the physical vector, $Q^{\prime}|\chi\rangle$ = 0, (for $|\chi\rangle=|\chi^{0}\rangle$) and from the set of reducible gauge transformations, $\delta|\chi\rangle$ = $Q^{\prime}|\chi^{1}\rangle$, $\ldots$, $\delta|\chi^{s-1}\rangle=Q^{\prime}|\chi^{s}\rangle$, for $gh(|\chi^{n}\rangle)=-n$, n=0,,sn=0,...,s, for some integer $s>0$ in the representation in $\mathcal{H}_{gh}$ in accordance with representation (2.1),
$\left(q_{i},p_{i},\eta_{12},\mathcal{P}_{12},\vartheta_{12},\lambda_{12},\mathcal{P}_{0}\right)|0\rangle$ $=$ $0,\;i=1,2$ (34)
it follows a finite sequence of relations underlying the $\eta^{i}_{G}$-independence of all of the above homogeneous in ghost number vectors:
$Q|\chi\rangle=0,$ $(\sigma^{i}+h_{i})|\chi\rangle=0,$ $\left(\varepsilon,{gh}_{H}\right)(|\chi\rangle)=(\varepsilon_{\chi},0),$ (35)
$\delta|\chi\rangle=Q|\chi^{1}\rangle,$ $(\sigma^{i}+h_{i})|\chi^{1}\rangle=0,$ $\left(\varepsilon,{gh}_{H}\right)(|\chi^{1}\rangle)=(\varepsilon_{\chi}+1,-1),$ (36)
$\ldots\ldots$ $\ldots\ldots$ $\ldots\ldots$
$\delta|\chi^{n-1}\rangle=Q|\chi^{n}\rangle,$ $(\sigma^{i}+h_{i})|\chi^{n}\rangle=0,$ $\left(\varepsilon,{gh}_{H}\right)(|\chi^{n}\rangle)=(\varepsilon_{\chi}+n\hskip-5.0pt\mod{}2,-n).$ (37)
By one of the peculiarity of BRST approach to mixed-antisymmetric HS fields is the fact that the Grassmann parity of any of $|\chi^{n}\rangle$ depends on the values of spin. The solutions of middle set of equations in (35)–(37) determine values of spin for the eigenvectors $\{|\chi^{n}\rangle_{{s[3]}}\}$ of generalized spin operator $(\sigma^{1},\sigma^{2},\sigma^{3})$ and possible values for parameters $h_{i}$:
$(h_{i}^{s[3]},h_{i}^{m|s[3]})$ $=$ $\left(-s_{i}+\textstyle\frac{d-6}{2}+2i\right)(1,1)+\left(0,{1}/{2}\right),\ \ i=1,2,3.$ (38)
As the result the operator $Q$ is nilpotent on the solutions for the middle set and the first equation in (35): $Q|\chi\rangle_{{s[3]}}=0$ is Lagrangian EoM and follows from gauge-invariant LF
$\mathcal{S}_{s[3]}=\int\hskip-1.99997ptd\eta_{0}\hskip-1.00006pt{}_{s[3]}\langle\chi|K_{s[3]}Q_{s[3]}|\chi\rangle_{s[3]}\hskip-1.00006pt\sim\hskip-1.00006pt\int\hskip-1.99997ptd^{d}x\hskip-1.49994pt\Big{(}\hskip-1.49994pt\Phi_{\mu[{s_{1}}],\nu[{s_{2}}],\rho[{s_{3}}]}\hskip-1.00006pt(\partial^{2}\hskip-1.00006pt+\hskip-1.00006ptm^{2})\hskip-1.00006pt\Phi^{\mu[{s_{1}}],\nu[{s_{2}}],\rho[{s_{3}}]}+\hskip-1.00006pt\texttt{more}\Big{)},$ (39)
(for substitution $(K_{s[3]},Q_{s[3]})=(K,Q)|_{h^{i}=h^{i}(s[3])}$) with reducible gauge transformations
$\delta|\chi^{0}\rangle_{s[3]}=Q_{s[3]}|\chi^{1}\rangle_{s[k]}\,,\ldots,\delta|\chi^{l}\rangle_{s[3]}=Q_{s[k]}|\chi^{l+1}\rangle_{s[3]},\ l=\sum_{i=1}^{3}s_{i}+2\,$, (40)
for independent gauge parameter $|\chi^{l+1}\rangle$. Here the usual inner product for the creation and annihilation operators is assumed with measure $d^{d}x$ over Minkowski space with additional terms: "more" with auxiliary fields differed for massless and massive basic HS field $\Phi^{\mu[{s_{1}}],\nu[{s_{2}}],\rho[{s_{3}}]}(x)$.
The LF appears by gauge theory for the particle with $(m,s[3])$ of $(\sum_{i=1}^{3}s_{i}+2)$ stage reducibility.
3 Lagrangian formulation with incomplete BRST operator \\
In this case the configuration space of fields contains less auxiliary fields as compared to one for LF with complete BRST operator. The LF is characterized by incomplete nilpotent, $Q_{c}$, BRST operator, operator of generalized spin $\sigma^{i}_{c}$ and BRST-extended operator holonomic constraints $\big{(}\widehat{L}_{ij},\widehat{T}_{ij}\big{)}$ from (27) not depending on conversion oscillators $b^{(+)}_{ij},d^{(+)}_{rs}$
$\sigma^{i}_{c}=\sigma^{i}_{c}=g_{0i}-q_{i}^{+}p_{i}-q_{i}p_{i}^{+},\quad\big{(}\widehat{L}_{ij},\widehat{T}_{ij}\big{)}=\big{(}\widehat{L}^{\prime}_{ij},\widehat{T}^{\prime}_{ij}\big{)}\big{|}_{(b^{(+)}_{ij},d^{(+)}_{rs})=0}$ (41)
The set of $Q_{c},\sigma^{i}_{c},\widehat{L}_{ij},\widehat{T}_{ij}$ acts in Hilbert space $\mathcal{H}^{f}\otimes H_{gh}^{o_{A}}$ and forms closed superalgebra/ It provides the consistent BRST $Q_{c}$-complex resolution to get LF with holonomic operator constraints imposed on whole set of restricted field $|\chi_{c}\rangle_{s[3]}$ and reducible gauge parameter $|\chi_{c}^{n}\rangle_{s[3]}$ vectors. The latter vectors have the decomposition in oscillators like (2.1) but without conversion $b^{(+)}_{ij},d^{(+)}_{rs}$ and ghost oscillators $\eta_{ij}^{+},\mathcal{P}_{ij}^{+},\vartheta_{rs}^{+},\lambda_{rs}^{+}$ for second-class constraint system.
The resulting LF for the particle with $(m,s[3])$ (using oscillators $f_{i},f_{i}^{+}$ for $m\neq 0$) after repeating the same steps as in the previous section has the form
$\mathcal{S}_{c|s[3]}(\chi_{c})=\int d\eta_{0}\;{}_{s[3]}\langle\chi_{c}|Q_{c}|\chi_{c}\rangle_{s[3]}\sim\hskip-1.00006pt\int\hskip-1.99997ptd^{d}x\hskip-1.49994pt\Big{(}\hskip-1.49994pt\Phi_{\mu[{s_{1}}],\nu[{s_{2}}],\rho[{s_{3}}]}\hskip-1.00006pt(\partial^{2}\hskip-1.00006pt+\hskip-1.00006ptm^{2})\hskip-1.00006pt\Phi^{\mu[{s_{1}}],\nu[{s_{2}}],\rho[{s_{3}}]}+\hskip-1.00006pt\texttt{more}\Big{)}\ ,$ (42)
$\Big{(}\delta;\ \widehat{L}_{ij},\ \widehat{T}_{rs}\Big{)}|\chi^{l}_{c}\rangle_{s[3]}=\Big{(}Q_{c}|\chi^{l+1}_{c}\rangle_{s[3]};\ {0},\ 0\Big{)},\ l=0,1,...,L=\sum_{i}s_{i}$ (43)
(with $\delta|\chi^{L+1}_{c}\rangle_{s[3]}=0$). For $\sum_{i}s_{i}=0$ (for scalar field) the theory appears by non-gauge one, whereas in general, it is the Abelian gauge model of $(\sum_{i}s_{i}-1)$-th stage reducibility.
In case of $\hat{s}_{3}=0$ LFs with (in)complete BRST operator corresponds to known [48] results for LFs for massless (massive) tensor fields $\Phi_{\mu^{1}[{\hat{s}_{1}}],\mu^{2}[{\hat{s}_{2}}]}$ subject to $Y[\hat{s}_{1},\hat{s}_{2}]$.
Note, the tensor expressions for the Lagrangian and tower of gauge transformations are easily obtained in calculating the inner products to reduce the oscillator presence. Second, the component form for the LF with incomplete BRST operator without off-shell holonomic constraints reveals the generalized gauge-invariant triplet formulation for reducible ISO(1,d1)ISO(1,d-1) group massless and massive representations with many spins $s^{\prime}[3]$ being fewer than $s[3]$. Third, according to the general result on equivalence for the dynamics for LFs with complete and incomplete BRST operator for irreducible mixed-symmetric HS fields $\Psi_{\mu^{1}({{s}_{1}}),\mu^{2}({{s}_{2}}),\mu^{3}({{s}_{3}})}$ subject to 3-row Young tableaux [57] the same principal statement follows for equivalence of dynamics generated by LFs with actions $\mathcal{S}_{s[3]}(\chi)$ and $\mathcal{S}_{c|s[3]}(\chi_{c})$ for the same field $\Phi_{\mu^{1}{[s_{1}]},\mu^{2}[{s_{2}}],\mu^{3}[{s_{3}}]}$ as well as with one determining by set of irrep conditions (3)–(6).
4 Deformation procedure for interacting vertices \\
To this end we introduce $p$, $p\geq 3$, copies of LFs (in adapting the model for Yang-Mills type interactions with gauge group $SU(N)$, for $p=N^{2}-1$) with vectors $|\chi^{(t)}_{c}\rangle_{{s}[k]_{t}}$, reducible gauge parameters $|\Lambda^{(t)n}_{c}\rangle_{{s}[k]_{t}}\equiv|\chi^{(t)n+1}_{c}\rangle_{{s}[k]_{t}}$, respective vacuum vectors $|0\rangle^{t}$ and oscillators for t=1,,pt=1,...,p (with notation $(s[3])_{t}\equiv(s[3]_{1},...,s[3]_{t})$ for arbitrary sets of indices for different copies). It permits one to determine the deformed action, gauge transformations up to $r$-tic vertices, r=3,4,,er=3,4,...,e in powers of deformation constant $g$, beginning from sum of $p$ LFs copies for free HS fields, then for cubic, quartic and so on vertices:
$S^{(m)_{p}}_{[e]|(s[3])_{p}}[(\chi_{c})_{p}]\ =\ \sum\nolimits_{t=1}^{p}\mathcal{S}_{c|s[3]_{t}}[\chi^{(t)}_{c}]+\sum\nolimits_{o=1}^{e}g^{o}S^{(m)_{p}}_{o|(s[3])_{p}}[(\chi_{c})_{p}],$ (44)
where
$\hskip-6.99997ptS^{(m)_{p}}_{1|(s[3])_{p}}[(\chi_{c})_{p}]=\sum_{1\leq i_{1}<i_{2}<i_{3}\leq p}\hskip-10.00002pt\int\prod_{j=1}^{3}d\eta^{(i_{j})}_{0}\Big{(}{}_{s[3]_{i_{j}}}\langle\chi^{(i_{j})}_{c}\big{|}V^{(3)}_{c}\rangle^{(m)_{(i)_{3}}}_{(s[3])_{(i)_{3}}}+h.c.\Big{)},$ (45)
$\hskip-8.99994ptS^{(m)_{p}}_{2|(s[3])_{p}}[(\chi_{c})_{p}]=\sum_{1\leq i_{1}<i_{2}<i_{3}<i_{4}\leq p}\hskip-10.00002pt\int\prod_{j=1}^{4}d\eta^{(i_{j})}_{0}\Big{(}{}_{s[3]_{i_{j}}}\langle\chi^{(i_{j})}_{c}\big{|}V^{(4)}_{c}\rangle^{(m)_{(i)_{4}}}_{(s[3])_{(i)_{4}}}+h.c.\Big{)},$ (46)
[…]
$\hskip-8.99994ptS^{(m)_{p}}_{e|(s[3])_{p}}[(\chi_{c})_{p}]=\sum_{1\leq i_{1}<i_{2}<...<i_{e}\leq p}\hskip-10.00002pt\int\prod_{j=1}^{e}d\eta^{(i_{j})}_{0}\Big{(}{}_{s[3]_{i_{j}}}\langle\chi^{(i_{j})}_{c}\big{|}V^{(e)}_{c}\rangle^{(m)_{(i)_{e}}}_{(s[3])_{(i)_{e}}}+h.c.\Big{)},$ (47)
also for deformed $l$-th level gauge transformations: $\delta^{l}_{[e]}|\Lambda^{(t)l}_{c}\rangle_{s[3]_{t}}=(\delta^{l}_{0}+\sum_{q=1}^{e}g^{q}\delta^{l}_{q})|\Lambda^{(t)l}_{c}\rangle_{s[3]_{t}}$ :
[…]
$\hskip-5.0pt\delta^{l}_{e}|\Lambda^{(t)l}_{c}\rangle_{{s}[3]_{t}}=-\sum_{1\leq i_{1}<...<i_{e-1}\leq p}\hskip-10.00002pt\int\prod_{j=1}^{e-1}d\eta^{(i_{j})}_{0}\Big{[}{}_{{s}[3]_{{i_{1}}}}\langle\chi^{(\{i_{1})}_{c}\big{|}\ldots{}_{{s}[3]_{{i_{e-2}}}}\langle\chi^{(i_{e-2})}_{c}\big{|}$ (49)
[…]
Here we have used the notations $(\chi_{c})_{p}=(\chi^{(1)}_{c},\chi^{(2)}_{c},...,\chi^{(p)}_{c})$, $|\Lambda^{(t)-1}_{c}\rangle\equiv|\chi^{(t)}_{c}\rangle$, the supersymmetrization of indices $\{i_{1},...,i_{e-1}\}$ with $r$-tic vertices $\big{|}{V}{}^{(r)l}_{c}\rangle$, for $l=-1,0,...,L_{t}$. The preservation for the interacting theory constructed from initial actions $\mathcal{S}^{m_{t}}_{0|\vec{s}_{k_{t}}}$, t=1,,pt=1,...,p the number $N_{t}$ of physical degrees of freedom (p.d.f.) determined by LFs for free HS field with spin ${s}[3]_{t}$, requires that the sum of all p.d.f. would be unchangeable, i.e. $\sum_{t}N_{t}=\mathrm{const}$. The property will be guaranteed, first, if the deformed action $S^{(m)_{p}}_{[e]|({s}[3])_{p}}$ will obey to sequence of new Noether identities in powers of $g$:
$g^{1}:$ $\delta_{0}S^{(m)_{p}}_{1|(s[3])_{p}}[(\chi_{c})_{p}]+\delta_{1}\mathcal{S}_{[0]|(s[3])_{p}}[(\chi_{c})_{p}]=0,$ (50)
$g^{2}:$ $\delta_{0}S^{(m)_{p}}_{2|(s[3])_{p}}[(\chi_{c})_{p}]+\delta_{1}S^{(m)_{p}}_{1|(s[3])_{p}}[(\chi_{c})_{p}]+\delta_{2}\mathcal{S}_{0|(s[3])_{p}}[(\chi_{c})_{p}]=0,$
$\ldots\ \ldots\ \ldots\ \ldots\ \ldots\ \ldots\$ $\ldots\ \ldots\ \ldots\ \ldots\$
$g^{e}:$ $\sum_{j=0}^{e}\delta_{j}S^{(m)_{p}}_{e-j|(s[3])_{p}}[(\chi_{c})_{p}]=0,$ (52)
second, from $\delta^{l}_{[\infty]}\delta^{l-1}_{[\infty]}|\Lambda^{(t)l-1}_{c}\rangle_{s[3]_{t}}|_{\partial S^{(m)_{p}}_{[\infty]|(s[3])_{p}}=0}=0$ for all levels of gauge transformations:
${g^{1}}:$ $\Big{(}\delta^{l}_{1}\delta^{l-1}_{0}+\delta^{{l}}_{0}\delta^{l-1}_{1}\Big{)}|\Lambda^{(t)l-1}_{c}\rangle_{\vec{s}_{k_{t}}}|_{\partial S^{(m)_{p}}_{[1]}=0}=0,$ (53)
${g^{2}}:$ $\Big{(}\delta^{l}_{2}\delta^{l-1}_{0}+\delta^{{l}}_{1}\delta^{l-1}_{1}+\delta^{l}_{0}\delta^{l-1}_{0}\Big{)}|\Lambda^{(t)l-1}_{c}\rangle_{s[3]_{t}}|_{\partial S^{(m)_{p}}_{[2]}=0}=0,$
[…]
$\vskip-2.15277pt{g^{e}}:$ $\Big{(}\delta^{l}_{e}\delta^{l-1}_{0}+\textstyle\sum_{p=1}^{e}\delta^{l}_{e-p}\delta^{l-1}_{p}\Big{)}|\Lambda^{(t)l-1}_{c}\rangle_{{s}[3]_{t}}|_{\partial S^{(m)_{p}}_{[e]}=0}=0$ (55)
(for $\delta^{-1}\equiv\delta$ and l=0,,L1l=0,...,L-1, where $L={\max}_{t}L_{(t)}$). Third, we should take into account for influence of traceless $L^{(t)}_{ij}$ and Young-symmetry $T^{(t)}_{rs}$ constraints on the structure of vertices $\big{|}V^{(q)}_{c}\rangle^{(m)_{(i)_{q}}}_{({s}[3])_{(i)_{q}}}$, $\big{|}{V}{}^{(q)l}_{c}\rangle^{(m)_{(i)_{q}}}_{({s}[3])_{(i)_{q}}}$ for q=3,4,,eq=3,4,...,e.
The resolution of (50) for xx=local cubic vertices with representation
$\big{|}V^{(3)}_{c}\rangle^{(m)_{(i)_{3}}}_{(s[3])_{(i)_{3}}}=\prod_{l=1}^{3}\delta^{(d)}\big{(}x-x_{i_{l}}\big{)}V^{(3)(m)_{(i)_{3}}}_{c|(s[3])_{(i)_{3}}}(x)\prod_{l=1}^{3}\eta^{(i_{l})}_{0}|0\rangle,\ \ |0\rangle\equiv\otimes_{t=1}^{p}|0\rangle^{t},$ (56)
which (e.g. given by Figure 1 for 2 massless and one massive HS fields) leads to the system
$\hskip-6.99997pt\mathcal{Q}(V^{(3)l-1}_{c|{(i)_{3}}},{V}^{(3)l}_{c|{(i)_{3}}})=\sum_{n=1}^{3}Q^{(i_{n})}\big{|}{V}^{(3)l-1}_{c}\rangle+Q^{(i_{t})}\Big{(}\big{|}V^{(3)l-1}_{c}\rangle-\big{|}{V}^{(3)l}_{c}\rangle\Big{)}=0$ (57)
(for $t=1,2,3$ and l=0,,L1l=0,...,L-1).
[…]
Figure 1: Cubic vertex for massive $\Phi^{(3)}_{\mu[s_{1}],\nu[s_{2}],\rho[s_{3}]}$ of spin $s[3]$ and two massless fields $\Phi^{(i)}_{\mu[{\lambda^{(i)}_{1}}],\nu[{\lambda^{(i)}_{2}}],\rho[{\lambda^{(i)}_{3}}]}$ of helicities $\lambda[3]_{i}$, $i=1,2$. The terms in $\ldots"$ correspond to the auxiliary fields from $|\chi^{(i)}\rangle_{s[3]_{i}}$
Its particular solution for coinciding $V^{(3)}_{c|{(i)_{3}}}={V}^{(3)0}_{c|{(i)_{3}}}=...={V}^{(3)k-1}_{c|{(i)_{3}}}$ has the usual form but augmented by the spin, traceless and Young-symmetry conditions with nilpotent total BRST operator $Q_{c}^{tot}$ in space $\otimes_{i=1}^{p}\mathcal{H}^{(i)}_{tot}$:
$Q_{c}^{tot}\big{|}V^{(3)}_{c}\rangle^{(m)_{(i)_{3}}}_{(s[3])_{(i)_{3}}}=0,\ Q_{c}^{tot}\equiv\sum_{i=1}^{p}Q_{c}^{(i)},\qquad\big{(}L^{(t)}_{ij},\,T^{(t)}_{rs}\big{)}\big{|}V^{(3)}_{c}\rangle^{(m)_{(i)_{3}}}_{({s}[3])_{(i)_{3}}}\ =\ 0,$ (58)
$\sigma^{(t)i_{t}}_{c}\big{|}V^{(3)}_{c}\rangle^{(m)_{(i)_{3}}}_{({s}[3])_{(i)_{3}}}\ =\ \Big{(}s_{i_{t}}+\frac{d-2+\theta_{m_{i},0}}{2}\Big{)}\big{|}V^{(3)}_{c}\rangle^{(m)_{(i)_{3}}}_{({s}[3])_{(i)_{3}}}.$ (59)
Without imposing algebraic constraints the deformed LF describes a dynamics for many particles with less spins then for original free HS fields $\Phi^{(p)}_{\mu[{\hat{s}{}^{(p)}_{1}}],\nu[{\hat{s}{}^{(p)}_{2}}],\rho[{\hat{s}{}^{(p)}_{3}}]}$. Thus, we developed the deformation procedure to get consistent interacting vertices with mixed-antisymmetric HS fields, starting from cubic ones to be explicitly considered in the subsequent research.
5 Conclusion
We developed BRST approach with complete and incomplete BRST operators to solve new problem of constructing LFs for mixed-antisymmetric HS field $\Phi_{\mu^{1}[{\hat{s}_{1}}],\mu^{2}[{\hat{s}_{2}}],\mu^{3}[{\hat{s}_{3}}]}$ of generalized integer spin $\mathbf{s}=(3,3,...,3,2,2,...,2,1,...,1)$ in a Minkowski space $\mathbb{R}^{1,d-1}$ whose symmetry of indices is subject to 3-column Young tableaux $Y[\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}]$. The respective LFs for free HS field have equivalent dynamics and are qualified as Abelian gauge theories of stages reducibility depending on sum of spin components equal respectively to $(\sum_{i}\hat{s}_{i}+2)$ and $(\sum_{i}\tilde{s}_{i}-1)$ with fewer auxiliary fields for the model with algebraic constraints. The results are applicable both for massless and massive particles. The Noether deformation procedure is proposed within approach with incomplete BRST operator for constructing interacting LF with local cubic vertices for $p$-copies of mixed-antisymmetric HS fields with preservation the Poincare group irreps for deformed (non-Abelian) gauge theory. The procedure permits to study interactions with totally-symmetric HS fields. The problem of constructing the BRST-BV (Batalin-Vilkovisky) minimal, quantum and effective actions for the interacting LF may be considered following to extension of BRST-BV approach [61] with incomplete, $Q_{c}$ BRST operator in [62] in order to perturbatively evaluate average expectation values of the quantities composed from mixed-antisymmetric HS fields.
Acknowledgements.
We wish to acknowledge the Organizing Committee of the XXV International Workshop-School QFTHEP’270, E. Boos for stimulating discussions and wonderful scientific spirit. A.R. is thankful to I. Buchbinder and to M. Vasiliev for useful comments.
CONFLICT OF INTEREST
The authors of this work declare that they have no conflicts of interest.